OpenStudy (anonymous):

(70000-x)*0.1+0.04x=4000 solve for x?

OpenStudy (tkhunny):

Have you considered first multiplying by 100?

OpenStudy (anonymous):

(70000*0.1)-(x*.1)+.04x=4000 7000-0.1x+0.04x=4000 7000-4000=0.1x-0.04x 3000=0.06x 3000/.06 = x 50000=x